OpenStudy (anonymous):

Find f(6) if f(x) = x 2 ÷ 3 + x. 4 10 18

OpenStudy (anonymous):

is that x^2?

OpenStudy (anonymous):

that is soppose to be a little 2 there isnt any of that ^

OpenStudy (anonymous):

\[x^{2}\]

OpenStudy (anonymous):

Yepp

OpenStudy (anonymous):

ok plug in 6 for x what do you get?

OpenStudy (anonymous):

6x i think?

OpenStudy (anonymous):

example f(a) f(x) = x + 2x - 3 f(a) = a + 2a -3

OpenStudy (anonymous):

ok im sorry i aint good at math at all lol

OpenStudy (anonymous):

you will be if you try

OpenStudy (anonymous):

yea i know i am trying

OpenStudy (anonymous):

just replace every x with 6

OpenStudy (anonymous):

okk

OpenStudy (anonymous):

the answer is 18

OpenStudy (anonymous):

correct